OpenStudy (fanduekisses):

What is the exact value of 20z over sqrt z^13 ?

OpenStudy (fanduekisses):

\[\left(\begin{matrix}20z \\ \sqrt{z ^{13}}\end{matrix}\right)\]

OpenStudy (fanduekisses):

That's the equation^

OpenStudy (anonymous):

20/z^5.5

OpenStudy (fanduekisses):

it cant be that these are the options...

OpenStudy (fanduekisses):

\[\left(\begin{matrix}20\sqrt{z} \\ z ^{5}\end{matrix}\right)\]

OpenStudy (anonymous):

20z/z^6.6=20/z^5.5 it is true

OpenStudy (anonymous):

no 20sq(z)/z^6

OpenStudy (fanduekisses):

yes thank you

OpenStudy (anonymous):

or z0/(z^5sq(z))

OpenStudy (anonymous):

20